# A Coherence-Based Measure of AGI

### v1

@@ 6 Conclusion
The arithmetic-mean definition of general intelligence, while simple and interpretable, conceals the structural imbalance of current AI systems. By generalizing this formulation to a continuum of compensability exponents, we reveal that apparent progress toward AGI is often confined to narrow strengths rather than coherent, system-wide competence. The proposed family of $AGI_{p}$ measures, and in particular the integrated $AGI_{\text{AUC}}$, provides a stricter and more interpretable framework for assessing generality.
Our results show that coherence—the balanced sufficiency of abilities across domains—is a more faithful indicator of general intelligence than arithmetic breadth alone. Genuine progress toward AGI will thus manifest not as higher averages, but as flatter and elevated $AGI_{p}$ curves reflecting resilient, interdependent capabilities. Incorporating coherence-aware metrics such as $AGI_{\text{AUC}}$ into future benchmarks can align empirical evaluation with the intended meaning of “general” in AGI.
@@ Appendix A Normalized Cognitive Capacity Tables (Percentage Form)
This appendix provides a detailed quantitative analysis of model performance across the ten cognitive domains defined in the A Definition of AGI framework [Hendrycks et al., 2025]. All subdomain values are drawn directly from Hendrycks et al. [2025] and are normalized to express the percentage of human-level proficiency attained within each subdomain. Subdomain weights follow the fixed allocations specified in that framework, reflecting each domain’s theoretical contribution to overall general intelligence; these weights may be refined in future extensions of the methodology.
Each subdomain carries a maximum weight that reflects its intended influence on the composite score. Subdomain raw results (drawb directly from [Hendrycks et al., 2025]) are normalized as:
@@ A.11 Discussion
AM captures domain breadth, WAM preserves the weight structure adopted by Hendrycks et al. [2025], while GM and WGM emphasize robustness and balance across subskills. Across Tables 312, the geometric aggregates consistently expose residual brittleness in multimodal perception, long-term memory (storage and retrieval), and adaptive reasoning—even where GPT-5’s AM/WAM approach human-level performance. This divergence between arithmetic and geometric summaries indicates that apparent gains are concentrated in a subset of faculties (e.g., crystallized knowledge and text-centric skills), with limited cross-modal integration and durable memory.
Methodologically, this paper uses the ten normalized domain scores reported by Hendrycks et al. [2025] as a shared empirical basis. We note, however, that these scores inherit several debatable design choices (e.g., potential conflation across subdomains and relatively low weights for adaptive planning). The contrast between AM/WAM and GM/WGM within Tables 312 further shows that arithmetic aggregation can inflate overall impressions of capability relative to coherence-oriented measures.
@@
In the main paper, we therefore focus on the aggregation principle rather than revising the underlying measurements. The reported domain scores are treated as reasonable, though potentially optimistic, proxies for current model ability, and are used to demonstrate how coherence-aware aggregation (culminating in $\mathrm{AGI}_{\mathrm{AUC}}$) provides a more faithful summary of systemic competence. Future work should revisit the domain taxonomy, explore nonlinear scaling and adaptive weightings based on empirical dependencies, and develop diagnostic tools to identify which deficiencies most constrain $\mathrm{AGI}_{\mathrm{AUC}}$—ultimately guiding targeted architectural or training interventions.

### v2

@@ 7 Conclusion
The arithmetic-mean definition of general intelligence, while simple and interpretable, conceals the structural imbalance of current AI systems. By generalizing this formulation to a continuum of compensability exponents, we reveal that apparent progress toward AGI is often confined to narrow strengths rather than coherent, system-wide competence. The proposed family of $\mathrm{AGI}_{p}$ measures, and in particular the integrated $\mathrm{AGI}_{\text{AUC}}$, provides a stricter and more interpretable framework for assessing generality.
Our results show that coherence, the balanced sufficiency of abilities across domains, is a more faithful indicator of general intelligence than arithmetic breadth alone. Genuine progress toward AGI will thus manifest not as higher averages, but as flatter and elevated $\mathrm{AGI}_{p}$ curves reflecting resilient, interdependent capabilities. Incorporating coherence-based metrics such as $\mathrm{AGI}_{\text{AUC}}$ into future benchmarks can align empirical evaluation with the intended meaning of “general” in AGI.
@@ Appendix A Normalized Cognitive Capacity Tables (Percentage Form)
This appendix provides a detailed quantitative analysis of model performance across the ten cognitive domains defined in the A Definition of AGI framework [hendrycks2025agidefinition]. All subdomain values are drawn directly from hendrycks2025agidefinition and are normalized to express the percentage of human-level proficiency attained within each subdomain. Subdomain weights follow the fixed allocations specified in that framework, reflecting each domain’s theoretical contribution to overall general intelligence; these weights may be refined in future extensions of the methodology.
Each subdomain carries a maximum weight that reflects its intended influence on the composite score. Subdomain raw results (drawb directly from [hendrycks2025agidefinition]) are normalized as:
@@ A.11 Discussion
AM captures domain breadth, WAM preserves the weight structure adopted by hendrycks2025agidefinition, while GM and WGM emphasize robustness and balance across subskills. Across Tables 312, the geometric aggregates consistently expose residual brittleness in multimodal perception, long-term memory (storage and retrieval), and adaptive reasoning—even where GPT-5’s AM/WAM approach human-level performance. This divergence between arithmetic and geometric summaries indicates that apparent gains are concentrated in a subset of faculties (e.g., crystallized knowledge and text-centric skills), with limited cross-modal integration and durable memory.
Methodologically, this paper uses the ten normalized domain scores reported by hendrycks2025agidefinition as a shared empirical basis. We note, however, that these scores inherit several debatable design choices (e.g., potential conflation across subdomains and relatively low weights for adaptive planning). The contrast between AM/WAM and GM/WGM within Tables 312 further shows that arithmetic aggregation can inflate overall impressions of capability relative to coherence-oriented measures.
@@
In the main paper, we therefore focus on the aggregation principle rather than revising the underlying measurements. The reported domain scores are treated as reasonable, though potentially optimistic, proxies for current model ability, and are used to demonstrate how coherence-aware aggregation (culminating in $\mathrm{AGI}_{\mathrm{AUC}}$) provides a more faithful summary of systemic competence. Future work should revisit the domain taxonomy, explore nonlinear scaling and adaptive weightings based on empirical dependencies, and develop diagnostic tools to identify which deficiencies most constrain $\mathrm{AGI}_{\mathrm{AUC}}$, ultimately guiding targeted architectural or training interventions.
Appendix B Beyond the CHC-domain Scores
Benchmark Description Gemini 3 Pro Gemini 2.5 Pro Claude Sonnet 4.5 GPT-5.1
Humanity’s Last Exam Academic reasoning 37.5 21.6 13.7 26.5
ARC-AGI-2 Visual reasoning puzzles 31.1 4.9 13.6 17.6
GPQA Diamond Scientific knowledge 91.9 86.4 83.4 88.1
AIME 2025 Mathematics (no tools) 95.0 88.0 87.0 94.0
MathArena Apex Challenging math contest problems 23.4 0.5 1.6 1.0
MMMU-Pro Multimodal understanding & reasoning 81.0 68.0 68.0 76.0
ScreenSpot-Pro Screen understanding 72.7 11.4 36.2 3.5
CharXiv Reasoning Information synthesis from charts 81.4 69.6 68.5 69.5
Video-MMMU Video-based knowledge acquisition 87.6 83.6 77.8 80.4
Terminal-Bench 2.0 Agentic terminal coding 54.2 32.6 42.8 47.6
SWE-Bench Verified Agentic coding (single attempt) 76.2 59.6 77.2 76.3
t2-bench Agentic tool use 85.4 54.9 84.7 80.2
FACTS Benchmark Suite Grounding, parametric, MM, search 70.5 63.4 50.4 50.8
SimpleQA Verified Parametric knowledge 72.1 54.5 29.3 34.9
MMLU Multilingual QA 91.8 89.5 89.1 91.0
Global PIQA Commonsense reasoning (100 languages) 93.4 91.5 90.1 90.9
MRCR v2 (8-needle) Long-context (128k avg) 77.0 58.0 47.1 61.6
Table 13: Comparison of leading large-language and multimodal models on diverse reasoning, multimodal, agentic, and long-context benchmarks. Scores are taken from the publicly released Gemini 3 Pro Model Evaluation Report [deepmind2025gemini3pro].
Figure 2: Comparison of model performance across aggregation exponents pp. Curves show $\mathrm{AGI}_{p}$ values derived from the 17 benchmarks in Table 13.
To illustrate that the coherence-based framework extends beyond the CHC taxonomy, we apply the same aggregation methodology to the 17 percentage-based benchmarks reported in Table 13, treating each as an independent capability axis. These benchmarks are drawn from the publicly released Gemini 3 Pro Model Evaluation Report [deepmind2025gemini3pro], which compares Gemini 3 Pro against GPT-5.1, Gemini 2.5 Pro, and Claude Sonnet 4.5 [claude_sonnet4_5].
These benchmarks span a broad spectrum of reasoning [phan2025humanity, chollet2025arc, rein2024gpqa, balunovic2025matharena], multimodal perception [yue2025mmmu, li2025screenspot, wang2024charxiv, hu2025video], agentic behavior [tbench_2025, yang2025swesmith, barres2025tau2], knowledge and grounding [haas2025simpleqa, chang2025global], and long-context understanding.
Let $b_{i}\in[0,100]$ denote a model’s score on benchmark $i$ for $i=1,\dots,17$. We compute $\mathrm{AGI}_{p}$ using the same continuous family of compensability exponents $p\in[-1,1]$ introduced in Eq. (2), with stability parameter $\varepsilon=10^{-6}$. We then evaluate the corresponding $\mathrm{AGI}_{p}$ curves and compute the integrated coherence score $\mathrm{AGI}_{\mathrm{AUC}}$ (Eq. 3) using the trapezoidal rule.
Results. The 17-benchmark evaluation produces a pattern that closely mirrors the CHC-domain analysis. Positive-$p$ aggregation (including the arithmetic mean at $p=1$) consistently overstates overall capability by masking unevenness across domains, whereas negative-$p$ aggregation sharply penalizes weak dimensions, revealing the true coherence profile of each model. Across the $p$-sweep, $\mathrm{AGI}_{\mathrm{AUC}}$ yields a more stable and discriminative ordering than the arithmetic mean: models with similar averages often diverge substantially in coherence, exposing persistent bottlenecks that single-number averages fail to capture.
Notably, Gemini 3 Pro exhibits the highest coherence across the full range of $p$ values. Its $\mathrm{AGI}_{p}$ curve remains consistently above those of the other models, reflecting improvements not only in overall performances but also in the hardest benchmarks, precisely the regime where coherence penalties matter most. This demonstrates that Gemini 3 Pro’s gains are broad rather than concentrated, leading to a substantially stronger AUC score (67%).
These findings confirm that the coherence-based evaluation introduced in this work is not dependent on the CHC taxonomy. Instead, it generalizes naturally to heterogeneous benchmark suites, offering a unified and domain-agnostic measure of holistic capability. By rewarding balanced competence and penalizing systemic weaknesses, $\mathrm{AGI}_{\mathrm{AUC}}$ provides a conservative and robust indicator of emerging general intelligence.
It is important to note that any aggregate score, whether derived from 17 benchmarks or hundreds, must be interpreted in light of the benchmark suite it summarizes. No numerical percentage can meaningfully reflect “progress toward AGI” unless the underlying tasks collectively span all essential cognitive faculties. When key abilities such as continual learning, long-term memory, or autonomous planning are missing, the resulting aggregate reflects performance only within the tested scope. A system with a bottleneck in any essential faculty cannot be described as achieving a given fraction of AGI, but rather as exhibiting unequal development across capabilities. Within any benchmark suite, however, coherence-based aggregation remains the most informative and principled evaluation method: it highlights imbalances, exposes hidden bottlenecks, and rewards the balance and robustness that are central to general intelligence.